OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -4).

OpenStudy (anonymous):

y =-1/4x^2 y2 = -4x y2 = -16x y =-1/16x^2

OpenStudy (mathmale):

You might want to abandon this account and choose a more suitable user name. Take a more positive approach! The standard form for the equation of a vertical parabolic graph (which this one is), with info on the location of the focus, is\[4py=x^2\] assuming that the vertex is at the origin. You are told that the focus is at (0,-4). Note that p is the distance from vertex to focus (or vice versa). What is p? Substitute that value of p into \[4py=x^2\] and simplify the resulting equation. Compare your result to the four possible answers.

OpenStudy (anonymous):

would p be 4?

OpenStudy (mathmale):

Yes. Continue with this problem solution, please.

OpenStudy (anonymous):

4(4)y=x^2, 16y=x^2 ?

OpenStudy (mathmale):

that's fine. Either leave your equation as is, or divide both sides by 16 to isolate y. Compare your result(s) to the four given answer choices.